OpenStudy (anonymous):

Find the fundamental matrix of the linear system x'=Ax where A is the 3x3 matrix [{1,1,0},{0,1,1},{0,0,1}. I know I should use the Putzer's Method but I keep getting stuck

OpenStudy (anonymous):

Finding the fundamental matrix is a matter of computing the eigenvalues and -vectors for the given matrix. Solve for all \(\lambda\): \[\begin{vmatrix} 1-\lambda&1&0\\ 0&1-\lambda&1\\ 0&0&1-\lambda \end{vmatrix}=0\]

OpenStudy (anonymous):

To get you started... With a cofactor expansion along the first column, you'll find the determinant to be \[\begin{vmatrix} \color{red}{1-\lambda}&1&0\\ \color{red}0&1-\lambda&1\\ \color{red}0&0&1-\lambda \end{vmatrix}=\color{red}{(1-\lambda)}\begin{vmatrix}1-\lambda&1\\0&1-\lambda\end{vmatrix}-\color{red}0+\color{red}0=(1-\lambda)^3\]

OpenStudy (anonymous):

right. I found that, but I was told to use the Putzer method since all of the eigenvalues are the same and will give the same eigenvectors

OpenStudy (anonymous):

I'm not familiar with the Putzer method, at least not by name, but to deal with repeated eigenvalues, the way I learned to do it was by finding the eigenvalues \(\vec{\eta}_i\) that satisfy \[(A-\lambda I)~\vec{\eta}_1=\vec{0}\\ (A-\lambda I)^2~\vec{\eta}_2=\vec{0}\\ (A-\lambda I)^3~\vec{\eta}_3=\vec{0}\] and I vaguely recall that you should get the same result if you do the following: \[(A-\lambda I)~\vec{\eta}_1=\vec{0}\\ (A-\lambda I)~\vec{\eta}_2=\vec{\eta}_1\\ (A-\lambda I)~\vec{\eta}_3=\vec{\eta}_2\] but I might be mistaken about that...

OpenStudy (anonymous):

hmmm I've never seen that but if it will help me solve the question I will try it. So am I trying to find when that matrix times the eigenvalue n will give 0 or is it when the determinant of that matrix times the eigenvalue is zero?

OpenStudy (anonymous):

You're multiplying matrices, not determinants. So the first eigenvalue can be found as follows: \[(A-I)~\vec{\eta}_1=\vec{0}~~\iff~~\begin{pmatrix}0&1&0\\0&0&1\\0&0&0\end{pmatrix}\begin{pmatrix}\eta_{1,1}\\\eta_{1,2}\\\eta_{1,3}\end{pmatrix}=\begin{pmatrix}0\\0\\0\end{pmatrix}\] for which you'll find that \(\eta_{1,1}\) can take on any number. For simplicity, we'd let \(\eta_{1,1}=1\), so the first eigenvector would be \[\vec{\eta}_1=\begin{pmatrix}1\\0\\0\end{pmatrix}\]

OpenStudy (anonymous):

this looks pretty similar to the putzer method. just different letters. I got stuck trying to figure out what n1, n2, and n3. I got n1=t^2/2(e^t)+1 n2=te^t snf n3=e^t-1 but I feel like these are wrong.

OpenStudy (anonymous):

By "n1,n2,n3", if you mean \(\vec{\eta}_1,~\vec{\eta}_2,~\vec{\eta}_3\), you should be getting vectors with constant entries... I'm not sure what you're saying you're getting. \[\begin{cases}(A-\lambda I)~\vec{\eta}_1=\vec{0}&(1)\\ (A-\lambda I)~\vec{\eta}_2=\vec{\eta}_1&(2)\\ (A-\lambda I)~\vec{\eta}_3=\vec{\eta}_2&(3)\end{cases}\] We already solved (1), so sub that into (2) to find \(\vec{\eta}_2\). \[\begin{pmatrix}0&1&0\\0&0&1\\0&0&0\end{pmatrix} \begin{pmatrix}\eta_{2,1}\\\eta_{2,2}\\\eta_{2,3}\end{pmatrix} =\begin{pmatrix}1\\0\\0\end{pmatrix}\] which gives \(\eta_{2,2}=1\) and \(\eta_{2,3}=0\), while \(\eta_{2,1}\) can take on any real number. Again, for simplicity, we can set \(\eta_{2,1}=1\), so the second eigenvector would be \[\vec{\eta}_2=\begin{pmatrix}1\\1\\0\end{pmatrix}\]

OpenStudy (anonymous):

oh i see where I messed up!!! thank you so much!!!

OpenStudy (anonymous):

You're welcome!